OpenStudy (anonymous):

Geometric sum sequence. Hello I'm working on a probability problem about two biased coins, but afterwards I arrive to a geometric series in my work that I am not sure how to calculate since I have not worked with series in a long time. I'll post my work below. will give a medal.

OpenStudy (anonymous):

\[\frac{p}{1-p}*\sum_{k=1}^{\infty}(1-p)^{3k}; 0<p<1\] attempt at the solution: know geometric series is \[\sum_{k=0}^{\infty} r^k=\frac{1}{1-r}\] so \[\sum_{k=1}^{\infty}r^{3k}=\frac{1}{1-r^3}-1?\] so my solution would be \[\frac{p}{1-p}(\frac{1}{1-(1-p)^3}-1)\] However I know this is not the solution, so I am not sure were I am making a mistake or if the formula I am using is correct. Any help is appreciated. Thank you!

OpenStudy (amistre64):

0 < p < 1 -1 < p-1 < 0 1 > 1-p > 0 well, since 1-p is less than 1, the series should converge using:\[\sum_{0}^{\infty}r^n=\frac{1}{1-r}\]

OpenStudy (amistre64):

are we sure that what we are starting with is valid to the overall solution?

OpenStudy (anonymous):

I know that the series converges, however I am not sure how to write what it converges to. Yes, it was a midterm question and I got taken off points for not arriving at the correct solution at the end.

OpenStudy (amistre64):

well, lets work this a longer way \[~~~~~~~S = r^3+r^6+r^9+...+r^{3k}\\ -r^3S = ~~~-r^6-r^9-...-r^{3k}-r^{3k+3}\\ -----------------\\ S=(r^3-r^{3k+3})/(1-r^3) \]

OpenStudy (amistre64):

simplifying we get: \[S=r^3\frac{1-r^{3k}}{1-r^3}\] as k to infinity, the top goes to 1 giving us:\[S=\frac{r^3}{1-r^3}\] seems fair to me

OpenStudy (anonymous):

May I know how you arrived to the part after you wrote"---------" Sorry my algebra is a little off and I don't see it.

OpenStudy (amistre64):

1S - r^3S = (1-r^3)S the right side of it is just subtracting all the common terms and we are left with r^3 - r^3(r^3k) we want to find S (the Sum) so we divide off the excess.

OpenStudy (anonymous):

I am not sure how it is 1S. Wouldn't an r^3 be missing in order to have a complete S value?

OpenStudy (anonymous):

I see you added it afterwards

OpenStudy (amistre64):

we want to find the sum of a set of terms: S = r^3 + r^6 + r^9 + ... + r^3k the historical approach to this is to subtract a modified version of itself from it to make life simpler: S - r^3 S r^3 S = r^6 + r^9 + ... + r^3k + r^3k r^3 so now all that we have to do is subtract one from the other S = r^3 + r^6 + r^9 + ... + r^3k -r^3 S = -r^6 - r^9 - ... - r^3k - r^3k r^3 ^^^^^^^^^^^^^^^^^ all these go to 0 therefore S - r^3 S = r^3 - r^3k r^3 ^^^^^^^ factor the S out S(1 - r^3) = r^3 - r^3k r^3 and divide off the excess to solve for S, the sum we wanted to know to begin with

OpenStudy (amistre64):

as k approaches infinity, r^3k goes to 0 since its less than 1, this leaves us with the formula: r^3/(1-r^3)

OpenStudy (anonymous):

As my solution I now get \[\frac{ p }{ 1-p }[\frac{(1-p)^3}{1-(1-p)^3}-1]\] However when I type in my answer to a site like wolfram alpha to compare my solutions don't match up: http://www.wolframalpha.com/input/?i=sum+from+1+to+infinity+of+p%281-p%29%5E%283k-1%29 http://www.wolframalpha.com/input/?i=p%2F%281-p%29%5B%281-p%29%5E3%2F%5B1-%281-p%29%5E3%5D-1%29%5D%5D Am I still making a mistake somewhere?

OpenStudy (amistre64):

that last -1 is not called for

OpenStudy (anonymous):

Why? Isn't it since I start from k=1 to infinity wouldn't I have to subtract the k=0 term?

OpenStudy (amistre64):

thats why i worked it the long way. S starts at r^3 and i simply took it from there to reach the finale mathing it out avoids trying to work out any memory tricks for adjustments

OpenStudy (amistre64):

the mathing process i worked out is correct: \[S=\frac{r^3}{1-r^3}\]

OpenStudy (amistre64):

if we had started with r^3(0) then we would had removed it, but i started at r^3(1)

OpenStudy (anonymous):

Oh I think I see it. So in order to work out something that is missing a few terms we subtract the initial terms missing? However I think my answer still does not match with the one on Wolfram Alpha. Is my substitution correct without the -1 at the end or did I make a small mistake still? Sorry for the plethora of questions.

OpenStudy (amistre64):

you have a few different wolf links, can you post the 'correct' one for me?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This one is the link to the solution they arrive at: http://www.wolframalpha.com/input/?i=sum+from+1+to+infinity+of+p%281-p%29%5E%283k-1%29 Afterwards I type in our answer and try to match it with theirs and seeing the variations and graphs to see if they match up. Here is the link to my answer: http://www.wolframalpha.com/input/?i=p%2F%281-p%29%5B%281-p%29%5E3%2F%5B1-%281-p%29%5E3%5D%29%5D%5D

OpenStudy (amistre64):

just a cursory working i get to \[\frac{p}{1-p}\frac{(1-p)(1-p)(1-p)}{1-(1-p)^3}\] \[\frac{p(1-p)(1-p)}{1-(1-p)^3}\] \[\frac{p(p^2+1-2p)}{1-(1-p)^3}\] \[\frac{p(p^2+1-2p)}{1-(1-3p+3p^2-p^3)}\] \[\frac{p(p^2+1-2p)}{3p-3p^2+p^3}\] \[\frac{p(p^2+1-2p)}{p(3-3p+p^2)}\] \[\frac{p^2-2p+1}{p^2-3p+3}\]

OpenStudy (anonymous):

I think it is almost the same solution since you can make the top \[(p-1)^3\] and include at the bottom (p-1) and factor out a negative one I see it now. Thank you so much. Maybe I mistyped it in wolfram alpha. I am not sure though

OpenStudy (amistre64):

sum of: p (1-p)^(3k-1) the wolf link to this is saying it doesnt understand your input

OpenStudy (anonymous):

I'll take a screen shot and attach it

OpenStudy (anonymous):

I'll retype it on wolfram alpha first. Sorry my screen shots came out dark gray

OpenStudy (amistre64):

your attempt is at trying:\[\frac{p}{1-p}~\frac{(1-p)^3}{\color{red}{1}}=p(1-p)^{3k-1}\] what happened to the demoninator of our summation?

OpenStudy (amistre64):

\[\sum(1-p)^{3k}=\frac{(1-p)^3}{1-(1-p)^3}\] \[\frac{p}{1-p}\sum(1-p)^{3k}=\frac{p}{1-p}~\frac{(1-p)^3}{1-(1-p)^3}\] oh wait, i see it: so you are trying:\[\sum p(1-p)^{3k-1}=\frac{p}{1-p}~\frac{(1-p)^3}{1-(1-p)^3} \]

OpenStudy (amistre64):

i say we test it out, let p=.3 :)

OpenStudy (anonymous):

Yes. Ok I will do that:)

OpenStudy (anonymous):

Yes it matched! Thank you so much. I understand it now :D Have a great day:D

OpenStudy (amistre64):

:) youre welcome